# Hierarchical World Models as Visual Whole-Body Humanoid Controllers

### v2

@@ 5 Related Work
Learning whole-body controllers for humanoids is a long-standing problem at the intersection of the machine learning and robotics communities. Humanoids are of particular interest to the learning community because of the high-dimensional nature of the problem (Heess et al., 2017; Merel et al., 2017; Peng et al., 2018; Merel et al., 2018b; Hasenclever et al., 2020; Wagener et al., 2022; Shi et al., 2022; Caggiano et al., 2022; Sferrazza et al., 2024), and to the robotics community because it is a promising morphology for general-purpose robotic agents (Grizzle et al., 2009; Li et al., 2023; BostonDynamics, 2024; Unitree, 2024; Cheng et al., 2024). Prior work predominantly focus on learning control policies for individual tasks using model-free reinforcement learning algorithms, many of which train behavior cloning (BC) policies (Wagener et al., 2022) and/or skill encoders (Heess et al., 2017; Merel et al., 2018b; Hasenclever et al., 2020; Shi et al., 2022) on human MoCap data (CMU, 2003) re-targeted to their considered humanoid embodiment. For example, Wagener et al. (2022) learn expert policies for each reference MoCap clip in the dataset, and then subsequently distills them into a single policy via supervised learning. Instead, we train a single tracking world model across all clips in the MoCapAct (Wagener et al., 2022) dataset, using only noisy rollouts and offline RL to do so. Concurrent work, HumanoidBench (Sferrazza et al., 2024), similarly introduces a whole-body control benchmark using the less expressive Unitree H1 (Unitree, 2024) embodiment. Our contributions differ in two important ways: (1) we develop a method for synthesizing natural human motions with a highly expressive humanoid model, while Sferrazza et al. (2024) focus on benchmarking existing methods for online RL, and (2) our proposed method additionally takes vision as input, whereas HumanoidBench considers tasks with privileged state information in their experiments.
@@
World models (and model-based RL more broadly) are of increasing interest to researchers due to their strong empirical performance in an online RL setting (Ha and Schmidhuber, 2018; Hafner et al., 2023; Hansen et al., 2024), as well as their promise of generalization to structurally similar problem instances (Zhang et al., 2018; Zheng et al., 2022; Lee et al., 2022; LeCun, 2022; Sobal et al., 2022; Brohan et al., 2023). Existing model-based RL algorithms can broadly be categorized into algorithms that select actions by planning with a learned world model (Ebert et al., 2018; Schrittwieser et al., 2020; Ye et al., 2021; SV et al., 2023; Hansen et al., 2024), and algorithms that instead learn a model-free policy using imagined rollouts from the world model (Kaiser et al., 2020; Hafner et al., 2023). We build upon the TD-MPC2 (Hansen et al., 2024) world model, which uses planning and has been shown to outperform other existing algorithms for continuous control (Hansen et al., 2023a; Lancaster et al., 2023; Feng et al., 2023; Sferrazza et al., 2024). In this work, we demonstrate that planning is key to success in the high-dimensional continuous control problems that we consider.
@@
Hierarchical RL offers a framework for subdividing a complex learning problem into more approachable subproblems, often by, e.g., leveraging (learned or manually designed) skill primitives (Pastor et al., 2009; Merel et al., 2017, 2018a; Shi et al., 2022) or facilitating learning over long time horizons via temporal abstractions (Nachum et al., 2019; LeCun, 2022; Hafner et al., 2022; Gumbsch et al., 2023; Chen et al., 2024). Our method, Puppeteer, is also hierarchical in nature, but does not rely on skill primitives nor temporal abstraction for task learning. Instead, we learn a single low-level world model that can be reused across a variety of downstream tasks, and instead introduce a hierarchy in terms of data sources and input modalities.
@@ 6 Conclusion
Puppeteer is a data-driven method for learning highly performant visual control policies for high-dimensional continuous control such as humanoids. We demonstrate that our method consistently produces motions that are considered natural and human-like by human evaluators compared to existing methods, and we find that our method is able to generalize zero-shot to gap lengths that far exceed the training data. These results are, to the best of our knowledge, unprecedented in the area of whole-body humanoid control. However, we acknowledge that our contributions have several limitations that may not be obvious:
  • While our proposed task suite consists of challenging visual whole-body control tasks with a detailed humanoid model, tasks primarily evaluate the visio-locomotive capabilities of current methods. However, humanoids hold the promise of performing highly dexterous tasks (such as, e.g., manipulating and carrying objects in unstructured environments alongside humans) which is not covered by our existing set of tasks. We expect development of new tasks will be increasingly important as algorithms continue to improve, and we hope that the release of our benchmark will help facilitate that.
  • Our hierarchical approach currently consists of two levels, only one of which is pretrained in the majority of our experiments. Our experiment on high-level pretraining suggests that it can be beneficial to pretrain both levels, but we expect this conclusion to vary depending on the choice of source and target tasks (Hansen et al., 2023b; Xu et al., 2023; Lin et al., 2023). Further research on how to pretrain and transfer the full hierarchical world model is warranted.
@@ Appendix A Additional Qualitative Results
Appendix B Implementation Details
@@
MoCap dataset. We use the “small” offline dataset provided by MoCapAct [Wagener et al., 2022], which is available at https://microsoft.github.io/MoCapAct. This dataset contains 20 noisy expert rollouts from each of 836 expert policies trained to track individual MoCap clips, totalling (suboptimal) 16,720 trajectories. Trajectories are variable length and are labelled with the CoMiC [Hasenclever et al., 2020] tracking reward which we use throughout this work.
@@
Puppeteer. We base our implementation off of TD-MPC2 and use default design choices and hyperparameters whenever possible. All hyperparameters are listed in Table 2. Our approach introduces only two new hyperparameters compared to prior work: loss coefficient for termination prediction (because our task suite has termination conditions; we add this to the TD-MPC2 baseline as well), and the number of low-level steps to take per high-level step (temporal abstraction).
@@
TD-MPC2. We use the official implementation available at https://github.com/nicklashansen/tdmpc2, but modify the implementation to support multi-modal observations and termination conditions as discussed in Section 3.
SAC. We benchmark against the implementation from https://github.com/denisyarats/pytorch_sac due to its strong performance on lower-dimensional DMControl tasks as well as its popularity among the community. We modify the implementation to support early termination.
DreamerV3. We use the official implementation available at https://github.com/danijar/dreamerv3, and use the default hyperparameters recommended for lower-dimensional DMControl tasks.
@@
Appendix C Task Suite
We propose a benchmark for visual whole-body humanoid control based on the “CMU Humanoid” model from DMControl [Tassa et al., 2018]. Our simulated humanoid has 56 fully controllable joints ($\mathcal{A}\in\mathbb{R}^{56}$), and includes both head, hands, and feet. Our task suite consists of 5 vision-conditioned whole-body locomotion tasks (corridor, hurdles, walls, gaps, stairs), as well as 3 tasks that use proprioceptive information only (stand, walk, run). All 8 tasks are illustrated in Figure 4.
Observations always include proprioceptive information, as well as either visual inputs (high-level agent) or a command (low-level agent). The proprioceptive state vector is $212$-dimensional and consists of relative joint positions and velocities, body velocimeter and accelerometer, gyro, joint torques, binary touch (contact) sensors, and orientation relative to world $z$-axis. Visual inputs are $64\times 64$ RGB images captured by a third-person camera (as seen in Figure 4), and commands are $15$-dimensional vectors with values in $[-1,1]$.
Appendix D User Study
To compare the “naturalness” of policies learned by our method vs. TD-MPC2, we design a user study in which humans are asked to watch short ($\sim$10101010s) pairs of clips of simulated humanoid motions generated for each of our 5 visual whole-body humanoid control tasks. Each user is presented with $2$ such pairs per task, totalling $10$ pairs per user. The concrete instructions provided to users are as follows:
@@ Appendix D User Study
In this study, you will watch pairs of short ( 10 seconds) clips of simulated humanoid motions. For each pair, you are asked to determine which of the two clips appear more "natural" and "human-like" to you, i.e., which clip looks more like the behavior of a real human.
@@
The order of clips is selected at random for each pair. Aggregate results from the user study are provided in Table 3, and Figure 10 shows a sample question from the user study. Participants are sourced from undergraduate and graduate student populations across multiple universities and disciplines. We do not collect personal or otherwise identifiable information about participants, and all participants have provided written consent to use of their responses for the purposes of this study.
@@
Table 3: Results from the user study. We summarize results from our user study ($n=46$) below by reporting per-pair aggregate numbers. Higher is better $\uparrow$. Clips generated by our method, Puppeteer, are considered more natural by a super-majority of participants.
@@
Pair 1 00 $2$ $44$
Pair 2 $1$ $3$ $42$
Aggregate $0.4\%$ $2.0\%$ $97.6\%$

### v3

@@ 5 Related Work
Learning whole-body controllers for humanoids is a long-standing problem at the intersection of the machine learning and robotics communities. Humanoids are of particular interest to the learning community because of the high-dimensional nature of the problem (Heess et al., 2017; Merel et al., 2017; Peng et al., 2018; Merel et al., 2018b; Hasenclever et al., 2020; Won et al., 2022; Wagener et al., 2022; Shi et al., 2022; Caggiano et al., 2022; Sferrazza et al., 2024; He et al., 2024; Cheng et al., 2024), and to the robotics community because it is a promising morphology for general-purpose robotic agents (Grizzle et al., 2009; Li et al., 2023; BostonDynamics, 2024; Unitree, 2024; Cheng et al., 2024). Prior work predominantly focus on learning control policies for individual tasks using model-free reinforcement learning algorithms, with human MoCap data (CMU, 2003) incorporated via either adversarial reward terms (Peng et al., 2018; 2021; 2022) or learned skill encoders (Heess et al., 2017; Merel et al., 2018b; Hasenclever et al., 2020; Shi et al., 2022; Wagener et al., 2022). While adversarial reward terms can produce natural behavior, this class of methods suffer from poor sample-efficiency as they learn a control policy from scratch for each downstream task. Our work is most similar to the latter class of methods, which enables reuse of the low-level policy and/or skill encoder across tasks. Most related to ours, MoCapAct (Wagener et al., 2022) first learn $\sim$2600 individual tracking policies via RL, then distill them into a multi-clip tracking policy via imitation learning, and subsequently train a high-level RL policy to output goal embeddings for the multi-clip policy to track. Their resulting representation is used to perform simple reaching and velocity control tasks from privileged state information in approx. $150$M environment steps. Our method trains a single world model to track the entire MoCap dataset, and is reused to learn a variety of visual whole-body control tasks in $\leq 3$M environment steps. Concurrent to our work, HumanoidBench (Sferrazza et al., 2024) similarly introduce a whole-body control benchmark using the less expressive Unitree H1 (Unitree, 2024) embodiment. Our contributions differ in two important ways: (1) we develop a method for synthesizing natural human motions with a highly expressive humanoid model while Sferrazza et al. (2024) benchmark existing methods for online RL without regard for naturalness, and (2) HumanoidBench solely considers tasks with privileged state information in their experiments (i.e., no visual observations).
@@
World models (and model-based RL more broadly) are of increasing interest to researchers due to their strong empirical performance in an online RL setting (Ha & Schmidhuber, 2018; Hafner et al., 2023; Hansen et al., 2024), as well as their promise of generalization to structurally similar problem instances (Zhang et al., 2018; Zheng et al., 2022; Lee et al., 2022; Xu et al., 2023; LeCun, 2022; Sobal et al., 2022; Brohan et al., 2023). Existing model-based RL algorithms can broadly be categorized into algorithms that select actions by planning with a learned world model (Ebert et al., 2018; Schrittwieser et al., 2020; Ye et al., 2021; SV et al., 2023; Hansen et al., 2024), and algorithms that instead learn a model-free policy using imagined rollouts from the world model (Kaiser et al., 2020; Hafner et al., 2023). We build upon the TD-MPC2 (Hansen et al., 2024) world model, which uses planning and has been shown to outperform existing algorithms for continuous control (Hansen et al., 2023a; Lancaster et al., 2023; Feng et al., 2023; Sferrazza et al., 2024). We demonstrate that planning is key to success in the high-dimensional continuous control problems that we consider.
@@
Hierarchical RL offers a framework for subdividing a complex learning problem into more approachable subproblems, often by, e.g., leveraging (learned or manually designed) skill primitives (Pastor et al., 2009; Merel et al., 2017; 2018a; Shi et al., 2022) or facilitating learning over long time horizons via temporal abstractions (Nachum et al., 2019; LeCun, 2022; Hafner et al., 2022; Gumbsch et al., 2023; Chen et al., 2024). Our method, Puppeteer, is also hierarchical in nature, but does not rely on skill primitives nor temporal abstraction for task learning. Instead, we learn a single low-level world model that can be reused across a variety of downstream tasks, and instead introduce a hierarchy in terms of data sources and input modalities.
@@ 6 Conclusion
We demonstrate that Puppeteer consistently produces motions that are considered natural and human-like by human evaluators compared to existing methods for visual RL, in an entirely data-driven manner and without any bells and whistles. These results are, to the best of our knowledge, unprecedented in the area of whole-body humanoid control. However, we acknowledge that our contributions have several limitations that may not be obvious: (i) while our proposed task suite consists of challenging visual whole-body control tasks with a detailed humanoid model, tasks primarily evaluate the visio-locomotive capabilities of current methods. We expect development of new tasks will be increasingly important as algorithms continue to improve, and we hope that the release of our benchmark will help facilitate that. (ii) Our hierarchical approach currently consists of two levels, only one of which is pretrained in the majority of our experiments. Our experiment on high-level pretraining suggests that it can be beneficial to pretrain both levels, but further research on how to pretrain and transfer the full hierarchical world model is warranted.
Ethics statement. All 51 participants in our user study are sourced from undergraduate and graduate student populations across multiple universities and disciplines on a volunteer basis. We do not collect personal or otherwise identifiable information about participants, and all participants have provided written consent to use of their responses for the purposes of this study. See Appendix F for more information.
Reproducibility statement. Code for method and environments, as well as model checkpoints, is made available at https://www.nicklashansen.com/rlpuppeteer. We rely on DMControl and MuJoCo for simulation which are publicly available and licensed under the Apache 2.0 license. We leverage the MoCapAct dataset for pretraining which is also publicly available and licensed under the MIT license. Implementation details and a full list of hyperparameters are provided in Appendix D.
Acknowledgements. This work was supported, in part, by NSF CCF-2112665 (TILOS). Nicklas Hansen is supported by NVIDIA Graduate Fellowship, and Vlad Sobal is supported by NSF Award 19922658.
@@ Appendix A Additional Qualitative Results
Appendix B Additional Quantitative Results
This section supplements our quantitative evaluations included in the main paper. In particular, Table 2 provides additional performance metrics for our low-level tracking agent, which we define as follows:
Success rate is defined as the percentage of clips that the tracking policy is able to track without ever exceeding a tracking error threshold of $0.5$ at any given time step. Because clips vary greatly in duration, we only evaluate success rate across at most 100 steps per clip by truncating longer clips.
Tracking error is defined as the mean distance between all joints and bodies in the humanoid wrt. the corresponding joints and bodies in the reference clip, evaluated on a per-time-step basis. Our definition of tracking error is consistent with that of Hasenclever et al. (2020) and Tassa et al. (2018).
CoMic score is the mean episode return achieved by the tracking policy as defined by the reward function proposed in CoMic (Hasenclever et al., 2020). We compute the CoMic score over all clips in the dataset, truncating clips to a maximum of 100 steps to prevent long clips from skewing results.
Table 2: Low-level tracking quality. We report three additional metrics of tracking quality for the low-level tracking agent: success rate as defined by Luo et al. (2023), average tracking error across all 836 clips, and CoMic (Hasenclever et al., 2020) reward across all clips.
Success rate (%) $\uparrow$ Tracking error $\downarrow$ CoMic score $\uparrow$
Offline only $6.2$ $0.503$ $42.6$
5% data $74.4$ $0.260$ $45.4$
25% data $79.6$ $0.225$ $46.3$
75% data $87.9$ $0.202$ $47.4$
Ours $\mathbf{88.3}$ $\mathbf{0.187}$ $\mathbf{48.7}$
Appendix C System Requirements
Detailed system requirements for our method and baselines are included in Table 3.
Table 3: System requirements. Training wall-time, inference time, and GPU memory requirements for Puppeteer, TD-MPC2, and SAC on a single NVIDIA GeForce RTX 3090 GPU. Overall, wall-time and system requirements of Puppeteer are mostly comparable to that of TD-MPC2 across both state-based and visual RL. SAC runs approx. 3.6x faster than Puppeteer, but does not achieve any meaningful performance. We exclude replay buffer memory requirements for clarity, but note that 1M transitions require 14.1 GB memory for visual RL and 1.0 GB memory for state-based RL. We store the replay buffer in GPU memory such that CPU and RAM usage is negligible.
Wall-time Inference time GPU memory
(h / 1M steps) (ms / step) (GB)
Puppeteer $21.8$ (vision: $29.0$) $88.2$ 0.6
TD-MPC2 $18.6$ (vision: $25.2$) $50.8$ 0.5
DreamerV3 $50.2$ $18.7$ 3.9
SAC $5.9$ $2.2$ 0.4
Appendix D Implementation Details
@@
MoCap dataset. We use the “small” offline dataset provided by MoCapAct (Wagener et al., 2022), which is available at https://microsoft.github.io/MoCapAct. This dataset contains 20 noisy expert rollouts from each of 836 expert policies trained to track individual MoCap clips, totalling (suboptimal) 16,720 trajectories. Trajectories are variable length and are labelled with the CoMiC (Hasenclever et al., 2020) tracking reward which we use throughout this work. We solely use this dataset during (pre)training of the low-level tracking agent; the high-level puppeteering agent is trained independently of the tracking agent using only online interaction data and task rewards.
@@
Puppeteer. We base our implementation off of TD-MPC2 and use default design choices and hyperparameters whenever possible. We experimented with alternative hyperparameters but did not observe any benefit in doing so. All hyperparameters are listed in Table 5. Our approach introduces only two new hyperparameters compared to prior work: loss coefficient for termination prediction (because our task suite has termination conditions; we add this to the TD-MPC2 baseline as well), and the number of low-level steps to take per high-level step (temporal abstraction).
@@
TD-MPC2. We use the official implementation available at https://github.com/nicklashansen/tdmpc2, but modify the implementation to support multi-modal observations and termination conditions as discussed in Section 3. We empirically observe that TD-MPC2 degenerates to highly unrealistic behavior without a contact-based termination condition.
SAC. We benchmark against the implementation from https://github.com/denisyarats/pytorch_sac (Yarats & Kostrikov, 2020) due to its strong performance on lower-dimensional DMControl tasks as well as its popularity among the community. We modify the implementation to support early termination. We experiment with a variety of design choices and hyperparameters as we find vanilla SAC to suffer from numerical instabilities on our task suite (presumably due to high-dimensional observation and action spaces), but are unable to achieve non-trivial performance. The ablation in Figure 8 (hierarchical planning) strongly suggests that planning is a key driver of performance in Puppeteer and TD-MPC2, while SAC is a model-free method incapable of planning. Design choices and hyperparameters that we experimented with are as follows:
Table 4: List of SAC design choices and hyperparameters. We experiment with a variety of design choices and hyperparameters, but find that they all fail to achieve non-trivial performance.
Design choice Values
Number of $Q$-functions $2,5$
TD-target Default, REDQ (Chen et al., 2021)
Activation ReLU, Mish, LayerNorm + Mish
MLP dim $256,512,1024$
Batch size $256,512$
Learning rate $3\times 10^{-4},1\times 10^{-3}$
DreamerV3. We use the official implementation available at https://github.com/danijar/dreamerv3, and use the default hyperparameters recommended for proprioceptive DMControl tasks. A key selling point of DreamerV3 is its robustness to hyperparameters across tasks (relative to SAC), but we find that DreamerV3 does not achieve any non-trivial performance on our task suite. While DreamerV3 is a model-based algorithm, it does not use planning, which the ablation in Figure 8 (hierarchical planning) finds to be a key driver of performance in Puppeteer and TD-MPC2.
@@
Appendix E Task Suite
We propose a benchmark for visual whole-body humanoid control based on the “CMU Humanoid” model from DMControl (Tassa et al., 2018). Our simulated humanoid has 56 fully controllable joints ($\mathcal{A}\in\mathbb{R}^{56}$), and includes both head, hands, and feet. Actions are normalized to be in $[-1,1]$. Our task suite consists of 5 vision-conditioned whole-body locomotion tasks (corridor, hurdles, walls, gaps, stairs), as well as 3 tasks that use proprioceptive information only (stand, walk, run). All 8 tasks are illustrated in Figure 4.
Observations always include proprioceptive information, as well as either visual inputs (high-level agent) or a command (low-level agent). The proprioceptive state vector is $212$-dimensional and consists of relative joint positions and velocities, body velocimeter and accelerometer, gyro, joint torques, binary touch (contact) sensors, and orientation relative to world $z$-axis. Visual inputs are raw $64\times 64$ RGB images captured by a third-person camera (as seen in Figure 4) without any preprocessing steps, and tracking commands are $15$-dimensional vectors (corresponding to $5$ points in 3D space) with values in $[-1,1]$.
Downstream task reward functions are based on the humanoid reward functions in DMControl with minimal modification to fit our higher DoF embodiment. All 5 visual tasks use the same reward function, which is proportional to forward velocity of the humanoid and is bounded to always be non-negative:
$R(\mathbf{s})\doteq\operatorname{clip}(\textnormal{linvel}_{x},[0,v_{\textnormal{target}}])$ (3)
where $\textnormal{linvel}_{x}$ is linear velocity along the $x$-axis, and the $\operatorname{clip}$ operator bounds the reward value to always be non-negative and at most that of a target velocity $v_{\textnormal{target}}$ which we set to $6$ in all tasks. The 3 proprioceptive tasks use a similar reward function, except that the agent is rewarded for velocity in any $XY$-direction, and has an additional term that encourages an upright pose:
$R(\mathbf{s})\doteq\operatorname{min}(|\textnormal{linvel}_{xy}|,v_{\textnormal{target}})+\alpha\cdot\textnormal{headpos}_{z}$ (4)
where $\alpha$ is a constant coefficient balancing the two reward terms, and $\textnormal{headpos}_{z}$ is the height of the humanoid head in the world frame. The additional height reward term is adopted from the stand, walk, and run run tasks that DMControl implement with a simplified humanoid model ($\mathcal{A}\in\mathbb{R}^{24}$). We find that the TD-MPC2 baseline produces very unrealistic behaviors without the additional reward term, so we choose to keep the term to make comparison more fair.
Appendix F User Study
To compare the “naturalness” of policies learned by our method vs. TD-MPC2, we design a user study in which humans are asked to watch short ($\sim$10101010s) pairs of clips of simulated humanoid motions generated for each of our 5 visual whole-body humanoid control tasks. Each user is presented with $2$ such pairs per task, totalling $10$ pairs per user. Sample clips used in the user study are available at https://www.nicklashansen.com/rlpuppeteer, as well as in Appendix A. Pairs are generated by converged Puppeteer and TD-MPC2 agents. We generate 5 rollouts per task for each of two separately trained agents (random seed $1$ and $2$) using the same method (i.e., Puppeteer or TD-MPC2), and select the clips with median episode return for each of the two random seeds. We use clips generated by two unique random seeds to ensure that diversity in behavior due to inter-seed variability is captured in the user study, and we select the median clip to ensure that we neither favor nor disadvantage a method due to outliers. The concrete instructions provided to users in the study are as follows:
@@ Appendix F User Study
In this study, you will watch pairs of short ($\sim$10 seconds) clips of simulated humanoid motions. For each pair, you are asked to determine which of the two clips appear more "natural" and "human-like" to you, i.e., which clip looks more like the behavior of a real human.
@@
The order of clips is selected at random for each pair. Aggregate results from the user study are provided in Table 6, and Figure 10 shows a sample question from the user study. Participants are sourced from undergraduate and graduate student populations across multiple universities and disciplines on a volunteer basis. We do not collect personal or otherwise identifiable information about participants, and all participants have provided written consent to use of their responses for the purposes of this study.
@@
Table 6: Results from the user study. We summarize results from our user study ($n=51$) below by reporting per-pair aggregate numbers. Higher is better $\uparrow$. Clips generated by our method, Puppeteer, are considered more natural by a super-majority of participants.
@@
Pair 1 00 $2$ $49$
Pair 2 $1$ $3$ $47$
Aggregate $0.4\%$ $1.8\%$ $97.8\%$